### v1

@@ 1 Introduction
A central conceptual shift inherent in this formulation is that entropy is a property of all states, not merely of stable equilibrium states. Stable equilibrium states (stable equilibrium state) are not the foundation of thermodynamics, but rather a distinguished subset of states characterized by extremal properties. Only for such states does entropy admit a fundamental equation whose derivatives define potentials such as temperature, pressure, and chemical potentials. Nonequilibrium states, by contrast, possess well-defined entropy values but no associated potentials — a distinction essential for both logical clarity and physical interpretation.
@@
A central point emphasized throughout this paper is not merely the applicability of thermodynamic concepts to small systems, but their universality: all concepts defined and employed here apply to all systems, independently of size, particle number, or degree of extensivity. No assumption of extensiveness is made at the foundational level. Properties such as entropy and energy are defined operationally for individual systems in individual states, rather than inferred from scaling arguments or ensemble limits. We purposely avoid to discuss here any instance or assumption of extensivity, because extensivity is not as a prerequisite for thermodynamic reasoning, but a contingent feature of certain classes of systems and states that, using the terminology introduced in (Gyftopoulos and Beretta, 1991, Ch. 16), can be modeled under the “simple-system model” approximation, also known as macroscopic limit. This emphasis is increasingly relevant in the context of what is often termed nanothermodynamics Hill (1964, 2001); Chamberlin (2015), where traditional macroscopic idealizations may fail, yet many thermodynamic results remain valid and indispensable.
@@ 3 The loaded meaning of the word “system”
The unit of measurement in the International System for the amount of a constituent is the “mole,” indicated by the symbol “mol” (sometimes also called the “gram-mole” and indicated by the symbol “gmol”), defined as the number of units (molecules, atoms, or particles) equal to the number of atoms in 12 grams of carbon-12. This number is called Avogadro’s number, $N_{\rm Av}=6.022\times 10^{23}$, and thus 1 mol (= 1 gmol) = $N_{\rm Av}$ particles. Of course, the International System prefixes for multiples are applicable. For example, for the kilomole, 1 kmol = $10^{3}$ mol.
@@ 5 What exactly do we mean by “state” of a system?
The fact that the list of values defining the state of a system at a given instant of time is infinite, [^5] In an engineering treatment, it will never be necessary to provide an explicit mathematical representation of the system’s state. However, it may be comforting to know that theoretical physics has developed the necessary tools. Just out of curiosity, if you are interested in further reading, we can mention that this infinite list of values is sometimes representable by a function, the wave function, but more generally, it must be represented by an operator, known as the “state or density operator,” acting on an appropriate Hilbert space associated with the system. means that a given system admits a vast multitude of possible states. Two identical systems are in two different states if the values of at least one property are different for the two systems: the two infinite lists of values defining the states of the two identical systems differ in at least one of the corresponding values.
@@ 6 Time evolution, interactions, and the concept of “process”
$\frac{\mathop{}\!\mathrm{d}A}{\mathop{}\!\mathrm{d}t}=f(A,{\rm internal\ forces},{\rm external\ forces})$ (2)
@@ 11 Notation for energy exchange and the energy balance equation
There are various types of equilibrium states. An unstable equilibrium is a state that can be induced to spontaneously evolve toward different often distant) state by some vanishingly small and brief interaction — an infinitesimal perturbation. Such a parturbation has only a temporary, negligible effect on the environment, leaving no permanent net external change. In contrast, a metastable equilibrium state is stable under small perturbations, but can changed without leaving permanent effects on the environment by means of larger perturbations.
@@ 16 Adiabatic availability: definition
Figure 8: Adiabatic availability $\Psi^{A}_{1}$ measures the maximum amount of energy that can be transferred out of system $A$ initially in state $A_{1}$ by means of a weight process. It obtains when the weight process for $A$ is reversible and the system ends in a stable equilibrium state. The final stable equilibrium state $A_{S_{1}}$ is uniquely determined by the initial state $A_{1}$.
@@ 18 Thermal reservoir: definition
That of thermal reservoir is a limiting concept, pedagogically useful to simplify practical modeling as well as theoretical development. But it is important to note that its defining condition provides only an approximate description of physical reality that holds with good accuracy for systems containing a large number of particles (exceedingly good for very large numbers, such as the triple-point model), but fails for systems with few particles. [^23] As a consequence, when a thermal reservoir $R$ is employed in the measurement procedure used to define the entropy difference between two states $A_{1}$ and $A_{2}$ of a system $A$ (see Sec. 21), the procedure requires measuring the energy change of $R$ in a reversible weight process of the composite system $AR$, during which the state of $A$ changes from $A_{1}$ to $A_{2}$. If system $A$ contains only a few particles while $R$ can only be realized with a large number of particles, the corresponding energy change of $R$ is exceedingly small and, in practice, difficult to detect. For this reason, in Ref. Beretta and Zanchini (2019) we developed a more elaborate — though pedagogically less practical — alternative formulation that avoids the use of thermal reservoirs as “entropymeters” and is better aligned with recent technological advances in the domain of small systems; see, e.g., Ref. Pekola and Karimi (2024). Moreover, its strict validity would constitute a violation of the second law. [^24] To see this, consider two identical reservoirs $R^{\prime}$ and $R^{\prime\prime}$, initially in the same state $R^{\prime}_{1}=R^{\prime\prime}_{1}$. Since they are both in MSE with $C$ in state $C_{R}$, they are in MSE with one another. Hence, state $R^{\prime}_{1}R^{\prime\prime}_{1}$ of the composite system $R^{\prime}R^{\prime\prime}$ is a stable equilibrium state with energy $E^{R^{\prime}R^{\prime\prime}}_{11}=E_{1}^{R^{\prime}}+E_{1}^{R^{\prime\prime}}$. Now, consider states $R^{\prime}_{2}$ and $R^{\prime\prime}_{2}$ such that $E_{2}^{R^{\prime}}=E_{1}^{R^{\prime}}+\Delta E$ and $E_{2}^{R^{\prime\prime}}=E_{1}^{R^{\prime\prime}}-\Delta E$. State $R^{\prime}_{2}R^{\prime\prime}_{2}$ is another stable equilibrium state for the composite system $R^{\prime}R^{\prime\prime}$. But $E^{R^{\prime}R^{\prime\prime}}_{22}=E^{R^{\prime}R^{\prime\prime}}_{11}$, so we would have two stable equilibrium states with the same values of energy, amounts, and parameters, in direct violation of the uniqueness asserted by the second law.
@@ 19 Available energy with respect to a thermal reservoir
Figure 10: Available energy $(\Omega^{R})_{1}^{A}$ with respect to thermal reservoir $R$ measures the maximum amount of energy that can be transferred out of system $A$ initially in state $A_{1}$ by means of a weight process for the composite system $AR$. It obtains independently of the initial stable equilibrium state $R_{1}$ of the reservoir when the weight process for $AR$ is reversible and the composite system $AR$ ends in a stable equilibrium state, i.e., AA and $R$ end in mutual stable equilibrium.
@@
A direct consequence of the results stated in this section, is that starting from any initial state $A_{1}$ of any system $A$ that can interact with a reservoir $R$ (initially in stable equilibrium state $R_{1}$), it is always possible, through a reversible weight process for the composite system $AR$ (see Figure 11), to reach a final state for $AR$ with values of $\boldsymbol{n}$} and $\boldsymbol{\beta}$} arbitrarily fixed among those compatible with the initial states and with system $A$ in an arbitrarily chosen final state $A_{2}$. In this case, in general, the system and the reservoir do not reach in mutual equilibrium. The final state $R_{2\text{rev}}$ of $R$ is uniquely determined by the chosen initial and final states $A_{1}$ and $A_{2}$ of the system, the initial state $R_{1}$, and the chosen values of $\boldsymbol{n}$} and $\boldsymbol{\beta}$}.
@@ 20 Temperature of a thermal reservoir: definition
We define the temperature of a reservoir, denoted as $T_{R}$, through the following measurement procedure, schetched in Figure 12.
@@ 23 Principle of entropy non-decrease in weight processes
with the condition that $(S_{\rm irr})_{12}=0$ if the weight process is reversible and $(S_{\rm irr})_{12}>0$ if the weight process is irreversible. During an irreversible weight process, the system loses some of its ability to transfer energy to a weight, and the entropy of the system increases. This increase, $(S_{\rm irr})_{12}$, is called entropy produced (or generated or created) in the system due to irreversibility. or simply entropy production (or generation or creation).
@@ 29 Temperature, total potentials, pressure
It can be shown (see below) that the equality of temperatures of two systems is a necessary condition for the two systems that can exchange energy to be in mutual equilibrium. The practical importance of this result arises from the fact that it allows for the indirect measurement of the temperature of a system $A$ (a partial derivative of its fundamental relation) by measuring the temperature of another system $B$ in mutual equilibrium with $A$. A thermometer is a system for which the temperature is easily measurable and the result readily displayed. By placing a thermometer $B$ in contact with system $A$ and waiting for mutual equilibrium to be reached, the temperature reading of the thermometer also provides the measurement of the temperature of $A$.
@@ 39 Clausius statement of the second law extended to nonequilibrium
To do this, consider again the interaction between $A$ and $B$ sketched in Figure 28 and suppose that we operate it as the result of the sequence of two separate processes with the assistance of a stationary or cyclic machine [^48] A stationary machine is a system that, while interacting with other systems, always remains in the same state. A cyclic machine is a system that, while interacting with other systems, periodically returns to its initial state. Note that the definition of a cyclic machine includes the case of a stationary machine. $X$ interposed between $A$ and $B$ as sketched in Figure 30. In the first process, machine $X$ receives from $A$ the amounts of energy $\delta E^{A\to X}$ and entropy $\delta S^{A\to X}$ respectively equal to the amounts $\delta E^{A\to B}$ and $\delta S^{A\to B}$ that in Figure 28 pass directly from $A$ to $B$. Machine $X$ uses them to attempt to separate part of the energy received from $A$ and to store it temporarily by lifting a weight $G$. System $A$ ends up in the same state $A_{2}$ as in Figure 28, while system $B$ ends up in a different state $B_{2^{\prime}}$. In the second process (Figure 31), the weight transfers to $B$ the energy it received from the machine, and $B$ ends up in the same state $B_{2}$ as in Figure 28. In the end, the sequence of two processes has the same net effects and, therefore, is equivalent to the direct exchange process in Figure 28. However, if the initial conditions of $A$ and $B$ are such that the amount of energy passed to the weight $G$ is non-negligible, we must conclude that a finite fraction of the energy received by $B$ (specifically, the energy it receives from the weight in the second process in Figure 31) is clearly identifiable as work. In this case, the non-work interaction between $A$ and $B$ is not heat.

### v2

@@ 1 Introduction
A central conceptual shift inherent in this formulation is that entropy is a property of all states, not merely of stable equilibrium states. Stable equilibrium states are not the foundation of thermodynamics, but rather a distinguished subset of states characterized by extremal properties. Only for such states does entropy admit a fundamental equation whose derivatives define potentials such as temperature, pressure, and chemical potentials. Nonequilibrium states, by contrast, possess well-defined entropy values but no associated potentials — a distinction essential for both logical clarity and physical interpretation.
@@
A central point emphasized throughout this paper is not merely the applicability of thermodynamic concepts to small systems, but their universality: all concepts defined and employed here apply to all systems, independently of size, particle number, or degree of extensivity. No assumption of extensiveness is made at the foundational level. Properties such as entropy and energy are defined operationally for individual systems in individual states, rather than inferred from scaling arguments or ensemble limits. We purposely avoid to discuss here any instance or assumption of extensivity, because extensivity is not a prerequisite for thermodynamic reasoning, but a contingent feature of certain classes of systems and states that, using the terminology introduced in (Gyftopoulos and Beretta, 1991, Ch. 16), can be modeled under the “simple-system model” approximation, also known as macroscopic limit. This emphasis is increasingly relevant in the context of what is often termed nanothermodynamics Hill (1964, 2001); Chamberlin (2015); Chamberlin and Lindsay (2024); Bedeaux et al. (2024), where traditional macroscopic idealizations may fail, yet many thermodynamic results remain valid and indispensable.
@@ 3 The loaded meaning of the word “system”
The unit of measurement in the International System for the amount of a constituent is the “mole,” indicated by the symbol “mol” (sometimes also called the “gram-mole” and indicated by the symbol “gmol”), defined as the number of units (molecules, atoms, or particles) equal to Avogadro’s number, $N_{\rm Av}=6.02214076\times 10^{23}$, and thus 1 mol (= 1 gmol) = $N_{\rm Av}$ particles. Of course, the International System prefixes for multiples are applicable. For example, for the kilomole, 1 kmol = $10^{3}$ mol.
@@ 5 What exactly do we mean by “state” of a system?
The fact that the list of values defining the state of a system at a given instant of time is infinite, means that a given system admits a vast multitude of possible states. Two identical systems are in two different states if the values of at least one property are different for the two systems: the two infinite lists of values defining the states of the two identical systems differ in at least one of the corresponding values.
@@
Representations of states across nonequilibrium thermodynamic frameworks
As noted in the Introduction, the profound significance of the laws of thermodynamics lies in their universal applicability across any level or framework of description chosen to model empirical reality. This holds provided the model possesses a fundamental mathematical structure and satisfies certain reasonable conditions — ensuring, for instance, that concepts like separability and statistical independence between the system and its surroundings are precisely defined.
Throughout two centuries of thermodynamic history, numerous frameworks have successfully modeled nonequilibrium phenomena. While these approaches vary significantly in their choice of independent properties defining the state space (in the sense of Eq. 1), most share common geometrical features within their mathematical structures, as captured for example by the different iterations of the GENERIC (General Equation for the Non-Equilibrium Reversible-Irreversible Coupling) construction Pavelka et al. (2018); Ottinger (2005); Grmela and Pavelka (2026) or variational formulations Gyarmati (1970); Ván and Kovács (2020).
The choice of independent properties used to define a “state” is not merely a matter of mathematical convenience; it reflects the physical scale, the degree of rarefaction, and the distance from equilibrium under consideration. Although these frameworks differ significantly in their mathematical formalisms, a common thread emerges: most modern approaches represent the state through various forms of probability distributions. Whether describing the likelihood of microstates in phase space, the distribution of internal mesoscopic configurations, or the statistical populations and coherences encoded in a density operator, these probabilistic representations provide a bridge between microscopic fluctuations and macroscopic observables.
The following list provides a minimal summary of the typical sets of independent properties adopted for state description in several prominent frameworks. The comparison highlights how different theories prioritize either a finite set of macroscopic fields or a more elaborate distribution function in order to capture the relevant physics. To ground these abstract descriptions in a concrete application, Section 51 presents an explicit example of a state formulation within the framework of quantum thermodynamics, illustrating how these principles manifest at the interface between information and energy.
  • Classical Statistical Mechanics Gibbs (1902); Landau and Lifshitz (1980): Probability distributions (densities) over classical phase space, representing the possible microstates defined by the positions and momenta of all particles.
  • Quantum Statistical Mechanics Tolman (1938); Pathria and Beale (2021): Probability distributions over a set of pure quantum states, typically expressed in terms of the eigenvectors of the system’s Hamiltonian operator.
  • Information-Theoretic Thermodynamics Jaynes (1957); Parrondo et al. (2015): Probability distributions over a discrete or continuous set of microstates or events, representing the observer’s uncertainty or informational description.
  • Stochastic Thermodynamics Seifert (2012); Sekimoto (2010); Risken (1996): Probability distributions over fluctuating trajectories and path-dependent variables (such as stochastic work and heat) that characterize the energetics of individual realizations.
  • Macroscopic Nonequilibrium Thermodynamics (CIT) Onsager (1931); De Groot and Mazur (1962); Ziegler (1983): Local-equilibrium field variables — functions of position and time — representing the macroscopic properties of continuum parcels assumed to be in locally stable equilibrium states.
  • Gradient (Non-local) Thermodynamics Cahn and Hilliard (1958); Antanovskii (1996); Ván (2003); Mauri (2013): A set of local field variables augmented by their spatial gradients, representing the state of non-uniform systems where the local energy density depends on the neighboring environment, such as in the description of diffuse interfaces and phase separation.
  • Continuum Mechanics Malvern (1969); Truesdell and Noll (2004); Gurtin et al. (2010): A finite set of local field variables, such as displacement, deformation gradient, and mass, momentum, and total energy densities, representing the macroscopic properties of continuum parcels under the local-equilibrium hypothesis.
  • Internal Variable Theories Berezovski and Ván (2017); Ván and Kovács (2020): A finite set of local field variables supplemented by hidden or internal variables that describe the microstructural state of a material (e.g., in thermoelasticity), often derived via variational principles to ensure thermodynamic consistency.
  • Extended Nonequilibrium Thermodynamics (EIT) Jou et al. (2010): A finite set of local field variables that includes classical densities augmented by dissipative fluxes (e.g., heat flux, viscous stress) treated as independent state variables in order to account for memory effects and finite signal propagation speeds away from equilibrium.
  • Mesoscopic Nonequilibrium Thermodynamics Mazur (1998); Rubi (2012); Anisimov and Longo (2024): Probability distributions over a set of local mesovariables describing internal configurations, such as the position of a Brownian particle, molecular orientation, cluster size, or degree of protein folding.
  • Rational Extended Thermodynamics (RET) Muller and Ruggeri (1998); Ruggeri and Sugiyama (2015): A finite set of local fields representing moments of the velocity distribution function, such as mass, momentum, and energy densities, as well as higher-order moments including heat flux and momentum flux.
  • Small-Scale and Rarefied Gas Dynamics Cercignani (1988); Struchtrup (2005): The local particle velocity distribution function, representing the probability of finding a molecule with a specified velocity at a given position and time, typically governed by the Boltzmann equation.
  • Chemical Kinetics Prigogine (1967); Beretta et al. (2012); Kondepudi and Prigogine (2014): A set of bulk or local macroscopic variables, such as species concentrations and reaction coordinates, sufficient to describe the chemical evolution of a reactive mixture toward stable equilibrium.
  • Quantum Thermodynamics Alicki and Lendi (2007); Deffner and Jarzynski (2013); Binder et al. (2018): The density operator, serving as a generalized probability distribution that represents both classical populations (probabilities) and quantum coherences.
  • Nanothermodynamics Hill (1964, 2001); Chamberlin (2015); Chamberlin and Lindsay (2024); Bedeaux et al. (2024): Statistical distributions over a large number of independent, small-scale replicas of a system, where the state is defined by finite-size parameters and internal degrees of freedom that account for non-extensive energy contributions and fluctuations.
While connecting these ideas to the extensive literature on nonequilibrium thermodynamics is valuable, acknowledging the many pioneers of the various approaches — and summarizing the different attempts at unification, construction — would require a level of discussion beyond the scope of this paper. Rather than claiming exhaustiveness, the preceding list is intended to illustrate the broad range of state representations to which the general thermodynamic concepts developed in this work may apply.
This diverse landscape of state representations illustrates the remarkable versatility of nonequilibrium thermodynamics across different physical scales and modeling priorities. It is important to emphasize that the foundational approach presented in this work is intended to be compatible with, and applicable within, any of these specific frameworks. By providing a rigorous set of operational definitions — grounded in the first and second laws as they apply to well-defined (separable and uncorrelated) systems — our construction establishes a self-consistent logical basis that helps clarify the limits of applicability of the assumptions underlying any chosen set of state variables. In particular, this grounding makes explicit the conditions of separability and absence of correlations required for the valid application of the energy balance and the principle of entropy non-decrease, thereby reducing the risk of conceptual misinterpretations when analyzing complex or interacting composite systems whose components become correlated during interaction.
@@ 6 Time evolution, interactions, and the concept of “process”
$\frac{\mathop{}\!\mathrm{d}A(t)}{\mathop{}\!\mathrm{d}t}=f(A(t),{\rm internal\ forces}(t),{\rm external\ forces}(t))$ (2)
@@ 11 Notation for energy exchange and the energy balance equation
There are various types of equilibrium states. An unstable equilibrium is a state that can be induced to spontaneously evolve toward different often distant) state by some vanishingly small and brief interaction — an infinitesimal perturbation. Such a perturbation has only a temporary, negligible effect on the environment, leaving no permanent net external change. In contrast, a metastable equilibrium state is stable under small perturbations, but can be changed without leaving permanent effects on the environment by means of larger perturbations.
@@ 16 Adiabatic availability: definition
Figure 8: Adiabatic availability $\Psi^{A}_{1}$ measures the maximum amount of energy that can be transferred out of system $A$ initially in state $A_{1}$ by means of a weight process. Such maximum is obtained when the weight process for $A$ is reversible and the system ends in a stable equilibrium state. The final stable equilibrium state $A_{S_{1}}$ is uniquely determined by the initial state $A_{1}$.
@@ 18 Thermal reservoir: definition
The concept of a thermal reservoir is an idealized limiting abstraction, providing a pedagogical framework that simplifies both practical modeling and theoretical development. But it is important to note that its defining condition provides only an approximate description of physical reality, that holds with good accuracy for systems containing a large number of particles (exceedingly good for very large numbers, such as the triple-point model), but fails for systems with few particles. [^22] As a consequence, when a thermal reservoir $R$ is employed in the measurement procedure used to define the entropy difference between two states $A_{1}$ and $A_{2}$ of a system $A$ (see Sec. 21), the procedure requires measuring the energy change of $R$ in a reversible weight process of the composite system $AR$, during which the state of $A$ changes from $A_{1}$ to $A_{2}$. If system $A$ contains only a few particles while $R$ can only be realized with a large number of particles, the corresponding energy change of $R$ is exceedingly small and, in practice, difficult to detect. For this reason, in Ref. Beretta and Zanchini (2019) we developed a more elaborate — though pedagogically less practical — alternative formulation that avoids the use of thermal reservoirs as “entropymeters” and is better aligned with recent technological advances in the domain of small systems; see, e.g., Ref. Pekola and Karimi (2024). Moreover, its strict validity would constitute a violation of the second law. [^23] To see this, consider two identical reservoirs $R^{\prime}$ and $R^{\prime\prime}$, initially in the same state $R^{\prime}_{1}=R^{\prime\prime}_{1}$. Since they are both in MSE with $C$ in state $C_{R}$, they are in MSE with one another. Hence, state $R^{\prime}_{1}R^{\prime\prime}_{1}$ of the composite system $R^{\prime}R^{\prime\prime}$ is a stable equilibrium state with energy $E^{R^{\prime}R^{\prime\prime}}_{11}=E_{1}^{R^{\prime}}+E_{1}^{R^{\prime\prime}}$. Now, consider states $R^{\prime}_{2}$ and $R^{\prime\prime}_{2}$ such that $E_{2}^{R^{\prime}}=E_{1}^{R^{\prime}}+\Delta E$ and $E_{2}^{R^{\prime\prime}}=E_{1}^{R^{\prime\prime}}-\Delta E$. State $R^{\prime}_{2}R^{\prime\prime}_{2}$ is another stable equilibrium state for the composite system $R^{\prime}R^{\prime\prime}$. But $E^{R^{\prime}R^{\prime\prime}}_{22}=E^{R^{\prime}R^{\prime\prime}}_{11}$, so we would have two stable equilibrium states with the same values of energy, amounts, and parameters, in direct violation of the uniqueness asserted by the second law.
@@ 19 Available energy with respect to a thermal reservoir
Figure 10: Available energy $(\Omega^{R})_{1}^{A}$ with respect to thermal reservoir $R$ measures the maximum amount of energy that can be transferred out of system $A$ initially in state $A_{1}$ by means of a weight process for the composite system $AR$. Such maximum is obtained independently of the initial stable equilibrium state $R_{1}$ of the reservoir when the weight process for $AR$ is reversible and the composite system $AR$ ends in a stable equilibrium state, i.e., AA and $R$ end in mutual stable equilibrium.
@@
A direct consequence of the results stated in this section, is that starting from any initial state $A_{1}$ of any system $A$ that can interact with a reservoir $R$ (initially in stable equilibrium state $R_{1}$), it is always possible, through a reversible weight process for the composite system $AR$ (see Figure 11), to reach a final state for $AR$ with values of $\boldsymbol{n}$} and $\boldsymbol{\beta}$} arbitrarily fixed among those compatible with the initial states and with system $A$ in an arbitrarily chosen final state $A_{2}$. In this case, in general, the system and the reservoir do not reach mutual equilibrium. The final state $R_{2\text{rev}}$ of $R$ is uniquely determined by the chosen initial and final states $A_{1}$ and $A_{2}$ of the system, the initial state $R_{1}$, and the chosen values of $\boldsymbol{n}$} and $\boldsymbol{\beta}$}.
@@ 20 Temperature of a thermal reservoir: definition
We define the temperature of a reservoir, denoted as $T_{R}$, through the following measurement procedure, sketched in Figure 12.
@@ 23 Principle of entropy non-decrease in weight processes
with the condition that $(S_{\rm irr})_{12}=0$ if the weight process is reversible and $(S_{\rm irr})_{12}>0$ if the weight process is irreversible. During an irreversible weight process, the system loses some of its ability to transfer energy to a weight, and the entropy of the system increases. This increase, $(S_{\rm irr})_{12}$, is called entropy produced (or generated or created) in the system due to irreversibility or simply entropy production (or generation or creation).
@@ 29 Temperature, total potentials, pressure
It can be shown (see below) that the equality of temperatures of two systems is a necessary condition for the two systems that can exchange energy to be in mutual equilibrium. The practical importance of this result arises from the fact that it allows for the indirect measurement of the temperature of a system $A$ (a partial derivative of its fundamental relation) by measuring the temperature of another system $B$ in mutual equilibrium with $A$. A thermometer is a system for which the temperature is easily measurable and the result readily displayed. By placing a thermometer $B$ in contact with system $A$ and waiting for mutual equilibrium to be reached, the temperature reading of the thermometer also provides the measurement of the temperature of AA. [^34] It should be noted that while a thermometer placed in contact with a system in a nonequilibrium state will provide a reading, this ‘apparent temperature’ is a property of the thermometer’s own stable equilibrium state after the system+thermometer composite reaches a specific steady-state or transient interaction, rather than an intrinsic property of the nonequilibrium system. In the present framework, we strictly distinguish between such interaction-dependent readings — which depend on the nature of the contact and the thermometer’s own properties — and the temperature, which is defined exclusively for stable equilibrium states via the entropy-energy relation.
@@ 39 Clausius statement of the second law extended to nonequilibrium
To do this, consider again the interaction between $A$ and $B$ sketched in Figure 28 and suppose that we operate it as the result of the sequence of two separate processes with the assistance of a stationary or cyclic machine [^48] A stationary machine is a system that, while interacting with other systems, always remains in the same state. A cyclic machine is a system that, while interacting with other systems, periodically returns to its initial state. Note that the definition of a cyclic machine includes the case of a stationary machine. $X$ interposed between $A$ and $B$ as sketched in Figure 30.494949It is important to emphasize that the cyclic machine introduced here is not an ‘engine’ in the sense of a practical application. Rather, it serves as a conceptual measuring device — a test system — that allows us to operationally determine if a non-work interaction can be decomposed into a work part and an irreducible non-work part. In the first process, machine $X$ receives from $A$ the amounts of energy $\delta E^{A\to X}$ and entropy $\delta S^{A\to X}$ respectively equal to the amounts $\delta E^{A\to B}$ and $\delta S^{A\to B}$ that in Figure 28 pass directly from $A$ to $B$. Machine $X$ uses them to attempt to separate part of the energy received from $A$ and to store it temporarily by lifting a weight $G$. System $A$ ends up in the same state $A_{2}$ as in Figure 28, while system $B$ ends up in a different state $B_{2^{\prime}}$. In the second process (Figure 31), the weight transfers to $B$ the energy it received from the machine, and $B$ ends up in the same state $B_{2}$ as in Figure 28. In the end, the sequence of two processes has the same net effects and, therefore, is equivalent to the direct exchange process in Figure 28. However, if the initial conditions of $A$ and $B$ are such that the amount of energy passed to the weight $G$ is non-negligible, we must conclude that a finite fraction of the energy received by $B$ (specifically, the energy it receives from the weight in the second process in Figure 31) is clearly identifiable as work. In this case, the non-work interaction between $A$ and $B$ is not heat.